lanaa:

or Solve using substitution. –7x + 9y = 12 x = –3

imagineeeeee:

I'm going to help you sense no one else is :/

imagineeeeee:

Substitute x=-3 into the equation so then you rewrite it as 7-(-3)+9y=12

imagineeeeee:

now you simplify that and it becomes 21+9y=12

imagineeeeee:

make the y go by itself so 9y=12-21 9y=-9

imagineeeeee:

then divide both sides of the equation by 9 y= -9/9 y=-1

imagineeeeee:

so the final part should be x=-3 and y=-1 or (-3,-1)

imagineeeeee:

hope that helpd you (:

lanaa:

thankss

imagineeeeee:

np (:

lanaa:

can u help me with some more if not its ok.

imagineeeeee:

yes I don't mind I'm not doing anything right now lol

imagineeeeee:

just send and I'll help (:

lanaa:

x = –3 –10x + 2y = 16

imagineeeeee:

what do you want me to do with this solve it?

lanaa:

yes its solve using substitution.

imagineeeeee:

okay do you remember what to do this is like really simple?

imagineeeeee:

so I'm gonna give you the break down for it okay?

imagineeeeee:

x= like any number you basically plug that into the equation and its connect to the first number and put in parenthesis like this -10x(-3)

imagineeeeee:

and so forth the rest of the equation

imagineeeeee:

next you simply the equation you know add what needs to be added and subtracted and multiplied and all that?

lanaa:

ok

imagineeeeee:

which would be 30+2y=16

imagineeeeee:

then you have to get y by its lonesome because y gets bullied

imagineeeeee:

then you dived both sides by which the both sides for this equation are y=-14/2 y= -7

imagineeeeee:

and then the final answer to them is x=-3 and y= -7 or (-3,-7)

imagineeeeee:

do you understand it a lil better?

lanaa:

yuh

imagineeeeee:

okay good

imagineeeeee:

glad I could help (:

notmeta:

is this algebra 1 or pre algebra?

Gucchi:

this easy asl no way u cant do this

notmeta:

@gucchi wrote:
this easy asl no way u cant do this
maybe its easy to you cause you've learned this years ago?

AcornTheNut:

exactly

ykbatman:

real.

ykbatman:

bro flexing knowing algerbra one while he in college

NaiNoah:

rizz?

MysticEcho:

real!!!!!!